Grade 12Physical Chemistry

5.6litgasat NTPweighs equal to 8g. the vapour- density of gas is?
  1. 32
  2. 16
  3. 8
  4. 40

Lavanya kunderu
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
Vapor thickness is the thickness of a vapor in connection to that of hydrogen. It might be characterized as mass of a specific volume of a substance partitioned by mass of same volume of hydrogen. 
vapor thickness = mass of n atoms of gas/mass of n particles of hydrogen. 
atomic mass of gas = 2*V.D.
8*22.4/(5.6) = 2 * V.D. 
Vapor DENSITY = 16